OpenStudy (anonymous):

What is the equation of a Parabola with the given vertex and focus vertex (-2,5) Focus (-2,6) Ayudame!

OpenStudy (anonymous):

Hallo?

Directrix (directrix):

Hi, have you done any work on this problem yet?

OpenStudy (anonymous):

I know it opens upwards, and has an axis of symmetry at -2. Not sure how to find the equation though

Directrix (directrix):

(x - h)^2 = 4*a*(y - 5) where (h,k) is the vertex, (-2,5) (x +2)^2 = 4*a* (y - 5)

Directrix (directrix):

Find a For veretical parabolawhere 'a' is the distance between vertex and focus.

OpenStudy (anonymous):

+1 right

Directrix (directrix):

a = 1

Directrix (directrix):

(y - 5) = (1/4) * (x + 2 )^2 In what form are you instructed to leave the equation?

OpenStudy (anonymous):

is the correct equation (x+2)^2+4(y-5) ?

OpenStudy (anonymous):

equation of a parabola is all that is says

OpenStudy (anonymous):

I'm assuming my answer was right, and the first equation you gave me is correct

Directrix (directrix):

(x+2)^2+4(y-5) ? Where is the equality symbol here.

Directrix (directrix):

@Phillip-boot

OpenStudy (anonymous):

meant =

OpenStudy (anonymous):

But that equation makes sense right?

Directrix (directrix):

Hold on while I think. Will you post the equation you wrote with the = symbol in it.

Directrix (directrix):

Click here and look at the graph: http://www.wolframalpha.com/input/?i=(y+-+5)+%3D+(1%2F4)+*+(x+%2B+2+)%5E2

OpenStudy (anonymous):

(x+2)^2=4(y-5)

OpenStudy (anonymous):

looks like it's the same graph

Directrix (directrix):

Our equations are equivalent. Same graph, yes. Vertex and Focus are correct.

OpenStudy (anonymous):

Killin it, thanks for checking/evaluating with me mate. You've gained a fan!

Directrix (directrix):

You are welcome. You have a new fan, too.